How many mol of zinc correspond to 3.0 × 102 atoms of zinc? Be sure to answer all parts. Enter your answer in scientific notation.

1
Expert's answer
2021-02-15T03:24:57-0500

Solution:

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Moles of zinc = 3.0×102 atoms × (1 mol / 6.022×1023 atoms) = 4.98×10-22 mol = 5.0×10-22 mol

Moles of zinc = 5.0×10-22 mol


Answer: 5.0×10-22 mol of zinc correspond to 3.0 × 102 atoms of zinc.